# Mining Your Own Secrets: Diffusion Classifier Scores for Continual Personalization of Text-to-Image Diffusion Models

### v2

@@ Abstract
Personalized text-to-image diffusion models have grown popular for their ability to efficiently acquire a new concept from user-defined text descriptions and a few images. However, in the real world, a user may wish to personalize a model on multiple concepts but one at a time, with no access to the data from previous concepts due to storage/privacy concerns. When faced with this continual learning (CL) setup, most personalization methods fail to find a balance between acquiring new concepts and retaining previous ones – a challenge that continual personalization (CP) aims to solve. Inspired by the successful CL methods that rely on class-specific information for regularization, we resort to the inherent class-conditioned density estimates, also known as diffusion classifier (DC) scores, for CP of text-to-image diffusion models. Namely, we propose using DC scores for regularizing the parameter-space and function-space of text-to-image diffusion models, to achieve continual personalization. Using several diverse evaluation setups, datasets, and metrics, we show that our proposed regularization-based CP methods outperform the state-of-the-art C-LoRA, and other baselines. Finally, by operating in the replay-free CL setup and on low-rank adapters, our method incurs zero storage and parameter overhead, respectively, over the state-of-the-art.
@@ 1 Introduction
Continual personalization (Smith et al., 2024b) aims to address the above challenge through continual learning (CL) and incorporating of new tasks with unseen concepts while retaining the previously acquired concepts. Naively adapting the existing (non-continual) personalization methods (Kumari et al., 2023; Gal et al., 2023) to acquire a new concept in CL setup often requires a complete retraining on data from all the seen concepts that the user desires to generate. However, storing a user’s personal data may raise resource/privacy concerns, and can be practically infeasible for edge devices. This calls for a replay-free CL solution, which does away with the seen data once a given concept has been acquired. C-LoRA (Smith et al., 2024b) handles these challenges by learning low-rank adapters (LoRAs) (Hu et al., 2022) per task, where each CL task acquires a single concept. It tackles the forgetting of previous concepts by penalizing the modification of any low-rank matrix spots allocated to these. We find that such a penalty leads to a degenerate CL solution where all LoRA parameter values are encouraged to be near zero (Sec. 3.1). This has further catastrophic consequences for the first task wherein the LoRA parameters are modified in an unconstrained manner and are thus more liable to the penalty. Nevertheless, given their edge in mitigating forgetting (Biderman et al., 2024), we resort to finetuning task-specific LoRAs while still consolidating their previous task knowledge for enabling CL. To the latter end, we note that merging the LoRA parameters based on task arithmetic (Ilharco et al., 2023) remains insufficient at retaining high generation quality. Instead, we propose to exploit the class-specific information of text-to-image diffusion models, with which we can consolidate the model’s discriminative semantic knowledge from previous tasks. Our inspiration for this comes from the broader CL literature on classification models where class-specific information, e.g., logits and softargmax scores, are often employed in countering forgetting with regularization (Li & Hoiem, 2017; Buzzega et al., 2020; Jha et al., 2024).
@@ 2 Background and Related Work
where $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\mathbf{c}$ is the conditioning information (e.g., class). Text-to-image diffusion (Saharia et al., 2022; Rombach et al., 2022) uses a cross-attention mechanism (Vaswani, 2017) in the U-Net (Ronneberger et al., 2015) to guide each step of the reverse process based on a text prompt $\mathbf{c}$. The keys $\mathbf{K}$ and values $\mathbf{V}$ to the cross-attention are rich semantic embeddings of $\mathbf{c}$ obtained from a pretrained text encoder $\phi$ (like CLIP (Radford et al., 2021a)).
@@
Continual Learning (CL) (Rolnick et al., 2019; Wang et al., 2024) aims to train a deep neural network on sequentially arriving tasks’ data to acquire new knowledge while retaining previously learned knowledge. A popular approach to CL comprises regularization-based methods that mitigate forgetting by imposing a penalty term on the learning objective. Based on how the penalty is computed, regularization may act on the parameter-space or the function-space. Parameter-space regularization, such as Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), constrains the changes to model weights that were important to previous tasks. EWC uses Fisher information (Fisher, 1922) to measure the parameter importance. Function-space regularization, like Learning without Forgetting (LwF) (Li & Hoiem, 2017) and Deep Model Consolidation (DMC) (Zhang et al., 2020), aims to preserve the model’s output behavior on previous tasks. These typically rely on knowledge distillation to ensure that the model’s predictions on previous tasks remain consistent. Our work uses EWC and DMC as representatives for the aforesaid regularization techniques.
@@
Classification with diffusion models (Li et al., 2023; Clark & Jaini, 2024) involves predicting how likely a class $\mathbf{c}^{i}$ is for an input $\mathbf{\mathbf{x}}$ by using a uniform Bayesian prior over all classes $\{\mathbf{c}^{1},\mathbf{c}^{2},..,\mathbf{c}^{N}\}$:
$p_{\theta}(\mathbf{\mathbf{c}}^{i}\mid\mathbf{\mathbf{x}})$ $=\frac{\exp\{-\mathbb{E}_{\mathbf{x},\epsilon,\mathbf{c}^{i},t}[\|\epsilon-\epsilon_{\theta}(\mathbf{\mathbf{x}}_{t},\mathbf{\mathbf{c}}^{i},t)\|^{2}]/\tau\}}{\sum_{j=1}^{N}\exp\{-\mathbb{E}_{\mathbf{x},\epsilon,\mathbf{c}^{j},t}[\|\epsilon-\epsilon_{\theta}(\mathbf{\mathbf{x}}_{t},\mathbf{\mathbf{c}}^{j},t)\|^{2}]/\tau\}},$ (3)
where $\tau>0$ is the temperature, and the probabilities $p_{\theta}=\{p_{\theta}(\mathbf{\mathbf{c}}^{1}\mid\mathbf{\mathbf{x}}),p_{\theta}(\mathbf{\mathbf{c}}^{2}\mid\mathbf{\mathbf{x}}),\ldots,p_{\theta}(\mathbf{\mathbf{c}}^{n}\mid\mathbf{\mathbf{x}})\}$ together comprise the Diffusion classifier (DC) scores. The expectation $\mathbb{E}$ is approximated over Monte-Carlo (MC) estimates across (inference) trials. Each trial samples a timestep $t\sim\mathcal{U}[1,T]$, computes a noisy input $x_{t}\sim q(\mathbf{x}_{t}|\mathbf{x}_{0})$ and then denoises it using the diffusion model $\epsilon_{\theta}$ conditioned on the class $\mathbf{c}^{i}$. DC scores thus help leverage a diffusion model’s rich pretrained generation knowledge for classification. Unlike existing works exploiting it only for zero-shot classification, we aim to use DC score as a regularization prior during training such that it can help mitigate forgetting in CL. We note that the computational costs for deriving DC scores are subject to the number of conditional inputs, and the number of trials. To circumvent these, existing works rely on iterative pruning of uninformative classes (Clark & Jaini, 2024), and appropriately choosing the diffusion timesteps across trials (Li et al., 2023). However, these methods still remain practically infeasible during training – iterative pruning per training iteration is computationally intensive while restricting the diffusion timesteps range leads to a loss in the signal reconstruction information. Accordingly, we propose practical considerations for efficient computation of DC scores during training.
@@ 3 Method
How do we structure our CL framework for DC scores? Using DC scores directly while acquiring new concepts can incur significant additional training cost (over single forward pass) given the need for several class-conditional forward passes per training image (Eq. 3). Instead, following Custom Diffusion (Kumari et al., 2023), we learn the $n^{\text{th}}$ concept with a new word vector $V_{n}^{*}$ and a LoRA layer by optimizing the diffusion loss (Eq. 1), the prior regularization loss using a common prior concept $\mathbf{c}^{0}$, and additionally a parameter-regularization loss in case of paremeter-space consolidation. After training, we freeze the word vector, and plug DC scores into two relatively shorter consolidation phases, one for each regularization method. Fig. 2 shows that these phases can work on their own as well as in tandem. Note that we train only one LoRA per task. After consolidation, the $n^{\text{th}}$ task LoRA serves two purposes: (a) handling inference-time queries for $\{1,2,\ldots,n\}$ tasks, (b) sequentially initializing the $(n+1)^{\text{th}}$ task LoRA. Next, we detail on each consolidation phase.
@@ 3.1 DC scores for Parameter-space Consolidation
$\mathcal{L}_{\text{ewc}}=\mathcal{L}_{\text{denoise}}+\delta\sum_{i=0}^{n}(2\mathbb{I}_{\{i=n\}}-1)\cdot H(p_{\theta},\mathbf{c}^{i}),$ (5)
where $\mathbb{I}$ is the indicator function, and $H(a,b)=-a\log b$ is the cross-entropy between the DC scores $p_{\theta}$ and the class label $\mathbf{c}^{i}$. The intuition behind Eq. 5 is that for the $n^{\text{th}}$ task images $\mathbf{x}_{n}$, the DC scores distribution ${p_{\theta}}$ should remain closer to the one-hot ground truth label $\mathbf{c}^{n}$, and should be further away from all other class labels $\mathbf{c}^{i<n}$. Fig. 3(a) and Algo. 2 outline our FIM computation framework.
@@
9:     $\mathbf{c}^{k}=\{\mathbf{c}^{0},\mathbf{c}^{j},\mathbf{c}^{n}\}$triangleright $\mathbf{c}^{0}=$ a prior concept
@@
11:     $\mathcal{L}_{\text{denoise}}=||\epsilon-\epsilon_{\theta_{s}}(\mathbf{x}_{t},\phi(\mathbf{c}^{n}),t)||^{2}_{2}$
@@
1:function GetDCS($n$, $\epsilon$, $\epsilon_{\theta}$, $t$, $\mathbf{x}_{t}$, $\mathbf{c}^{k}$)
2:Input: $n$: number of seen concepts, $\epsilon$: noise, $\epsilon_{\theta}$: diffusion model, $t$: timestep, $x_{t}$: noisy input, $\mathbf{c}^{k}$: a subset of seen concepts to be considered for computing DC scores
@@
4:     $\mathsf{pred}[\mathbf{c}^{i}]=\epsilon_{\theta}(\mathbf{x}_{t},\phi(\mathbf{c}^{i}),t)$ for $\mathbf{c}^{i}\in\mathbf{c}^{k}$
5:     $\mathcal{L}_{\text{denoise}}[\mathbf{c}^{i}]=||\epsilon-\mathsf{pred}[\mathbf{c}^{i}]||^{2}_{2}$ for $\mathbf{c}^{i}\in\mathbf{c}^{k}$
@@
7:     ${p_{\theta}}$ = $\mathbf{c}^{i}$: $\omega$ for $i\in[0,n]$} $\triangleright$ DC scores placeholder
8:     ${p_{\theta}}[\mathbf{c}^{i}]=\frac{\exp\{-\mathcal{L}_{\text{denoise}}[\mathbf{c}^{i}]/\tau\}}{\sum_{j=1}^{|\mathbf{c}^{k}|}\exp\{-\mathcal{L}_{\text{denoise}}[\mathbf{c}^{j}]/\tau\}}$ for $\mathbf{c}^{i}\in\mathbf{c}^{k}$
@@
3:     $\mathbf{c}^{k-2}=\{\mathbf{c}^{i}\sim[\mathbf{c}^{1},\mathbf{c}^{n-1}]\}$triangleright sample w/o replacement
4:     $\mathbf{c}^{k}=\mathbf{c}^{k-2}$+$\{\mathbf{c}^{0},\mathbf{c}^{n}\}$triangleright $\mathbf{c}^{0}=$ a prior concept
@@
6:     $\mathcal{L}_{\text{denoise}},{p_{\theta}},\text{\textendash}$ – = GetDCS($n$, $\epsilon$, $\epsilon_{\theta_{n}}$, $t$, $\mathbf{x}_{t}$, $\mathbf{c}^{k}$)
@@
11:        $\mathcal{L}_{\text{ewc}}\mathrel{+}=\mathcal{L}_{\text{denoise}}[\mathbf{c}^{i}]+\delta\mathcal{L}_{\text{DC}}$
@@
Large number of seen concepts: Note that the number of class-conditional forward passes for DC score derivation grows linearly with the number of concepts. Here, iteratively pruning the uninformative classes (Clark & Jaini, 2024) still requires multiple passes. Instead, we propose reducing the cost of forward passes to a constant factor using a subset $\mathbf{c}^{k}$ of the number of seen concepts for DC scores computation. This subset always comprises at least two concepts: the task-shared prior concept $\mathbf{c}^{0}$, and the ground truth current task’s concept $\mathbf{c}^{n}$. On top of these, we randomly sample $|k-2|$ previous concepts without replacement from the set $\{\mathbf{c}^{1},...,\mathbf{c}^{n-1}\}$, where $k>2$ is a hyperparameter chosen by grid search. Note that freezing the word vector $V^{*}_{i}$ helps us compute the textual embedding per concept $\phi(\mathbf{c}^{i})$ once and reuse it throughout the consolidation phase.

### v3

@@ Abstract
Personalized text-to-image diffusion models have grown popular for their ability to efficiently acquire a new concept from user-defined text descriptions and a few images. However, in the real world, a user may wish to personalize a model on multiple concepts but one at a time, with no access to the data from previous concepts due to storage/privacy concerns. When faced with this continual learning (CL) setup, most personalization methods fail to find a balance between acquiring new concepts and retaining previous ones – a challenge that continual personalization (CP) aims to solve. Inspired by the successful CL methods that rely on class-specific information for regularization, we resort to the inherent class-conditioned density estimates, also known as diffusion classifier (DC) scores, for CP of text-to-image diffusion models. Namely, we propose using DC scores for regularizing the parameter-space and function-space of text-to-image diffusion models, to achieve continual personalization. Using several diverse evaluation setups, datasets, and metrics, we show that our proposed regularization-based CP methods outperform the state-of-the-art C-LoRA, and other baselines. Finally, by operating in the replay-free CL setup and on low-rank adapters, our method incurs zero storage and parameter overhead, respectively, over C-LoRA. Project page.
@@ 1 Introduction
Continual personalization (Smith et al., 2024b) aims to address the above challenge through continual learning (CL) and incorporating of new tasks with unseen concepts while retaining the previously acquired concepts. Naively adapting the existing (non-continual) personalization methods (Kumari et al., 2023; Gal et al., 2023) to acquire a new concept in CL setup often requires a complete retraining on data from all the seen concepts that the user desires to generate. However, storing a user’s personal data may raise resource/privacy concerns, and can be practically infeasible for edge devices. This calls for a replay-free CL solution, which does away with the seen data once a given concept has been acquired. C-LoRA (Smith et al., 2024b) handles these challenges by learning low-rank adapters (LoRAs) (Hu et al., 2022) per task, where each CL task acquires a single concept. It tackles the forgetting of previous concepts by penalizing the modification of any low-rank matrix spots allocated to these. We find that such a penalty leads to a degenerate CL solution where all LoRA parameter values are encouraged to be near zero (Sec. 3.1). This has further catastrophic consequences for the first task wherein the LoRA parameters are modified in an unconstrained manner and are thus more liable to the penalty. Nevertheless, given their edge in mitigating forgetting (Biderman et al., 2024), we resort to finetuning task-specific LoRAs while still consolidating their previous task knowledge for enabling CL. To the latter end, we note that merging the LoRA parameters based on task arithmetic (Ilharco et al., 2023) remains insufficient at retaining high generation quality. Instead, we propose to exploit the class-specific information of text-to-image diffusion models, with which we can consolidate the model’s discriminative semantic knowledge from previous tasks. Our inspiration for this comes from the broader CL literature on classification models where class-specific information, e.g., logits and softargmax scores, are often employed in countering forgetting with regularization (Li & Hoiem, 2017; Buzzega et al., 2020; Jha et al., 2023; 2024).
@@ 2 Background and Related Work
where $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\mathbf{c}$ is the conditioning information (e.g., class). Text-to-image diffusion (Saharia et al., 2022; Rombach et al., 2022) uses a cross-attention mechanism (Vaswani, 2017) in the U-Net (Ronneberger et al., 2015) to guide each reverse process step with a text prompt encompassing $\mathbf{c}$. The keys $\mathbf{K}$ and values $\mathbf{V}$ to the cross-attention are rich semantic embeddings of $\mathbf{c}$ obtained from a pretrained text encoder $\phi$ (like CLIP (Radford et al., 2021a)).
@@
Continual Learning (CL) (Rolnick et al., 2019; Jha et al., 2020) aims to train a deep neural network on sequentially arriving tasks’ data to acquire new knowledge while retaining previously learned knowledge. A popular approach to CL comprises regularization-based methods that mitigate forgetting by imposing a penalty term on the learning objective. Based on how the penalty is computed, regularization may act on the parameter-space or the function-space. Parameter-space regularization, such as Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), constrains the changes to model weights that were important to previous tasks. EWC uses Fisher information (Fisher, 1922) to measure the parameter importance. Function-space regularization, like Learning without Forgetting (LwF) (Li & Hoiem, 2017) and Deep Model Consolidation (DMC) (Zhang et al., 2020), aims to preserve the model’s output behavior on previous tasks. These typically rely on knowledge distillation to ensure that the model’s predictions on previous tasks remain consistent. Our work uses EWC and DMC as representatives for the aforesaid regularization techniques.
@@
Classification with diffusion models (Li et al., 2023; Clark & Jaini, 2024) involves predicting how likely a class $\mathbf{c}_{i}$ is for an input $\mathbf{x}$ by using a uniform Bayesian prior over all classes $\{\mathbf{c}_{1},\mathbf{c}_{2},..,\mathbf{c}_{N}\}$:
$p_{\theta}(\mathbf{c}_{i}\mid\mathbf{x})$ $=\frac{\exp\{-\mathbb{E}_{\mathbf{x},\epsilon,\mathbf{c}_{i},t}[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{c}_{i},t)\|^{2}]/\tau\}}{\sum_{j=1}^{N}\exp\{-\mathbb{E}_{\mathbf{x},\epsilon,\mathbf{c}_{j},t}[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{c}_{j},t)\|^{2}]/\tau\}},$ (3)
where $\tau>0$ is the temperature, and the probabilities $p_{\theta}=\{p_{\theta}(\mathbf{c}_{1}\mid\mathbf{x}),p_{\theta}(\mathbf{c}_{2}\mid\mathbf{x}),\ldots,p_{\theta}(\mathbf{c}_{n}\mid\mathbf{x})\}$ together comprise the Diffusion classifier (DC) scores. The expectation $\mathbb{E}$ is approximated over Monte-Carlo (MC) estimates across (inference) trials. Each trial samples a timestep $t\sim\mathcal{U}[1,T]$, computes a noisy input $x_{t}\sim q(\mathbf{x}_{t}|\mathbf{x}_{0})$ and then denoises it using the diffusion model $\epsilon_{\theta}$ conditioned on the class $\mathbf{c}_{i}$. DC scores thus help leverage a diffusion model’s rich pretrained generation knowledge for classification. Unlike existing works exploiting it only for zero-shot classification, we aim to use DC score as a regularization prior during training such that it can help mitigate forgetting in CL. We note that the computational costs for deriving DC scores are subject to the number of conditional inputs, and the number of trials. To circumvent these, existing works rely on iterative pruning of uninformative classes (Clark & Jaini, 2024), and appropriately choosing the diffusion timesteps across trials (Li et al., 2023). However, these methods still remain practically infeasible during training – iterative pruning per training iteration is computationally intensive while restricting the diffusion timesteps range leads to a loss in the signal reconstruction information. Accordingly, we propose practical considerations for efficient computation of DC scores during training.
@@ 3 Method
How do we structure our CL framework for DC scores? Using DC scores directly while acquiring new concepts can incur significant additional training cost (over single forward pass) given the need for several class-conditional forward passes per training image (Eq. 3). Instead, following Custom Diffusion (Kumari et al., 2023), we learn the $n^{\text{th}}$ concept with a new word vector $V_{n}^{*}$ and a LoRA layer by optimizing the diffusion loss (Eq. 1), the prior regularization loss using a common prior concept $\mathbf{c}_{0}$, and additionally a parameter-regularization loss in case of paremeter-space consolidation. After training, we freeze the word vector, and plug DC scores into two relatively shorter consolidation phases, one for each regularization method. Fig. 2 shows that these phases can work on their own as well as in tandem. Note that we train only one LoRA per task. After consolidation, the $n^{\text{th}}$ task LoRA serves two purposes: (a) handling inference-time queries for $\{1,2,\ldots,n\}$ tasks, (b) sequentially initializing the $(n+1)^{\text{th}}$ task LoRA. Next, we detail on each consolidation phase.
@@ 3.1 DC scores for Parameter-space Consolidation
$\mathcal{L}_{\text{ewc}}=\mathcal{L}_{\text{denoise}}+\delta\mathbb{E}_{\mathbf{x}}\Big{[}\sum_{i=0}^{n}(2\mathbb{I}_{\{i=n\}}-1)\cdot H(p_{\theta},\mathbf{c}_{i})\Big{]},$ (5)
where $\mathbb{I}$ is the indicator function, and $H(a,b)=-a\log b$ is the cross-entropy between the DC scores $p_{\theta}$ and the class label $\mathbf{c}_{i}$. The intuition behind Eq. 5 is that for the $n^{\text{th}}$ task images $\mathbf{x}_{n}$, the DC scores distribution ${p_{\theta}}$ should remain closer to the one-hot ground truth for concept $\mathbf{c}_{n}$, and should be farther from all other class labels $\mathbf{c}_{i<n}$. Fig. 3(a) and Algo. 2 outline our FIM computation framework.
@@
9:     $\mathbf{c}_{k}=\{\mathbf{c}_{0},\mathbf{c}_{j},\mathbf{c}_{n}\}$triangleright $\mathbf{c}_{0}=$ a prior concept
@@
11:     $\mathcal{L}_{\text{denoise}}=||\epsilon-\epsilon_{\theta_{s}}(\mathbf{x}_{t},\phi(\mathbf{c}_{n}),t)||^{2}_{2}$
@@
1:function GetDCS($n$, $\epsilon$, $\epsilon_{\theta}$, $t$, $\mathbf{x}_{t}$, $\mathbf{c}_{k}$)
2:Input: $n$: number of seen concepts, $\epsilon$: noise, $\epsilon_{\theta}$: diffusion model, $t$: timestep, $x_{t}$: noisy input, $\mathbf{c}_{k}$: a subset of seen concepts to be considered for computing DC scores
@@
4:     $\mathsf{pred}[\mathbf{c}_{i}]=\epsilon_{\theta}(\mathbf{x}_{t},\phi(\mathbf{c}_{i}),t)$ for $\mathbf{c}_{i}\in\mathbf{c}_{k}$
5:     $\mathcal{L}_{\text{denoise}}[\mathbf{c}_{i}]=||\epsilon-\mathsf{pred}[\mathbf{c}_{i}]||^{2}_{2}$ for $\mathbf{c}_{i}\in\mathbf{c}_{k}$
@@
7:     ${p_{\theta}}$ = $\mathbf{c}_{i}$: $\omega$ for $i\in[0,n]$} $\triangleright$ DC scores placeholder
8:     ${p_{\theta}}[\mathbf{c}_{i}]=\frac{\exp\{-\mathcal{L}_{\text{denoise}}[\mathbf{c}_{i}]/\tau\}}{\sum_{j=1}^{|\mathbf{c}_{k}|}\exp\{-\mathcal{L}_{\text{denoise}}[\mathbf{c}_{j}]/\tau\}}$ for $\mathbf{c}_{i}\in\mathbf{c}_{k}$
@@
3:     $\mathbf{c}_{k-2}=\{\mathbf{c}_{i}\sim[\mathbf{c}_{1},\mathbf{c}_{n-1}]\}$triangleright sample w/o replacement
4:     $\mathbf{c}_{k}=\mathbf{c}_{k-2}$+$\{\mathbf{c}_{0},\mathbf{c}_{n}\}$triangleright $\mathbf{c}_{0}=$ a prior concept
@@
6:     $\mathcal{L}_{\text{denoise}},{p_{\theta}},\text{\textendash}$ – = GetDCS($n$, $\epsilon$, $\epsilon_{\theta_{n}}$, $t$, $\mathbf{x}_{t}$, $\mathbf{c}_{k}$)
@@
11:        $\mathcal{L}_{\text{ewc}}\mathrel{+}=\mathcal{L}_{\text{denoise}}[\mathbf{c}_{i}]+\delta\mathcal{L}_{\text{DC}}$
@@
Large number of seen concepts: Note that the number of class-conditional forward passes for DC score derivation grows linearly with the number of concepts. Here, iteratively pruning the uninformative classes (Clark & Jaini, 2024) still requires multiple passes. Instead, we propose reducing the cost of forward passes to a constant factor using a subset $\mathbf{c}_{k}$ of the number of seen concepts for DC scores computation. This subset always comprises at least two concepts: the task-shared prior concept $\mathbf{c}_{0}$, and the ground truth current task’s concept $\mathbf{c}_{n}$. On top of these, we randomly sample $|k-2|$ previous concepts without replacement from the set $\{\mathbf{c}_{1},...,\mathbf{c}_{n-1}\}$, where $k>2$ is a hyperparameter chosen by grid search. Note that freezing the word vector $V^{*}_{i}$ helps us compute the textual embedding per concept $\phi(\mathbf{c}_{i})$ once and reuse it throughout the consolidation phase.